A potential difference of 100 V produces a current of 3.0 A in a certain resistor. (a) What is the resistance of the resistor? (b) What is the current when the potential difference is 25 V?

Please show work.
Answer #1

a) By Ohm's law

V = IR

R = V/I = 100 / 3 = 33.33 ohm

b) I = V/R

= 25 / 33.33

= 0.75 A